### v3

# A Non-Relativistic Limit for Heterotic Supergravity and its Gauge Lagrangian

Abstract
We show that the $D=10$ heterotic supergravity in a non-relativistic (NR) limit has a finite Lagrangian due to non-trivial cancellations of divergent parts arising from the Chern-Simons terms in the curvature of the $\hat{B}$-field and the Yang-Mills Lagrangian. This is similar to what happens in bosonic supergravity between the Ricci scalar, $\hat{R}$, and the $-\frac{1}{12}\hat{H}^{2}$ term, under the same limit. In this work we present the explicit form of the gauge transformations and curvatures after considering a NR limit and we compute the full finite Lagrangian in its covariant form. Particularly, we address both the ungauged part and the gauged part, and we write it using covariant derivatives and curvatures for the gauge fields. These results constitutes the first computation of a finite Lagrangian for heterotic supergravity under these kinds of limits. As an interesting feature of the expansion implemented, the Green-Schwarz mechanism for the two-form can be trivialized and terms equivalent to Chern-Simons contributions naturally arise from the previous property.
1 Introduction
In the last decade there has been huge progress made in the study of the different formulations of non-relativistic string theory [1]-[3] and its supergravity limit [4]-[17]  [^1] See [18]-[20] for reviews on this topic.. The low energy limit of bosonic string theory is given by the universal NS-NS supergravity, whose fundamental fields $\hat{g}$, $\hat{B}$, $\hat{\phi}$ admit a non-relativistic expansion of the form [13],
$\hat{g}_{\mu\nu}$ $=$ $c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\eta_{ab}+h_{\mu\nu}\,,$ (1.1)
$\hat{g}^{\mu\nu}$ $=$ $\frac{1}{c^{2}}{\tau^{\mu}}_{a}{\tau^{\nu}}_{b}\eta^{ab}+h^{\mu\nu}\,,$ (1.2)
$\hat{B}_{\mu\nu}$ $=$ $-c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\epsilon_{ab}+b_{\mu\nu}\,,$ (1.3)
$\hat{\phi}$ $=$ $\ln(c)+\varphi\,,$ (1.4)
where $\mu,\nu=1,\dots,10$ are space-time indices and we have split the flat index $\hat{a}=(a,a^{\prime})$ where $a,b,...=0,1$ are the transverse directions and $a^{\prime},b^{\prime},...=2,\dots,9$. The quantities in the previous expansion obey the following Newton-Cartan relations,
$\tau_{\mu}{}^{a}e^{\mu}{}_{a^{\prime}}$ $=\tau^{\mu}{}_{a}e_{\mu}{}^{a^{\prime}}=0\,,\qquad e_{\mu}{}^{a^{\prime}}e^{\mu}{}_{b^{\prime}}=\delta^{a^{\prime}}_{b^{\prime}}\,,$ (1.5)
$\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (1.6)
While, apriori, this expansion produces divergent tensorial quantities, i.e., Riemann tensor and its traces, H strength, etc., the NS-NS supergravity Lagrangian,
$S=\int d^{10}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}\Big{(}\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\Big{)}\,,$ (1.7)
with
$\hat{H}_{\mu\nu\rho}=3\partial_{[\mu}\hat{B}_{\nu\rho]}\,,$ (1.8)
remains finite due to a curious cancellation between the divergent part of the Ricci scalar and the divergent part of the $\hat{H}^{2}$ term [13].
In the case of the bosonic sector of heterotic supergravity, we expect a similar cancellation when including the gauge field $\hat{A}_{\mu}^{i}$, $i=1,\dots,N$ and $N=496$ which is the dimension of the heterotic gauge group. Given a particular expansion for the gauge field, the inclusion of Chern-Simons terms in the curvature of the $\hat{B}$-field should produce divergences that must be compensated by the Yang-Mills Lagrangian. While there are several candidates for the expansion of the A-field in the heterotic scenario [21]-[22], there is still no explicit computation of the finite form of the heterotic Lagrangian under this limit. The main reason of an explicit computation of this Lagrangian is that exact results under the NR limit usually contain large contributions when they are not written in covariant form. We expect that the NR limit of the heterotic supergravity admits a rewriting in terms of $O(D,D+N)$ multiplets. T-duality is an exact symmetry of string theory and the low energy limit of the theory should be able to written in terms of a generalized metric and dilaton. In other words, we expect that the formalism of Double Field Theory [23]-[26] could be able to rewrite the theory, when the fundamental degrees of freedom are parametrized using a non-Riemannian ansatz [27]-[30].
On the other hand, we expect that the vielbein remains gauge invariant while the $\hat{B}$-field inherits a non-covariant gauge transformation (Green-Schwarz mechanism), similarly to what happens in the relativistic case, i.e.,
$\delta_{\hat{\lambda}}\hat{A}_{\mu}^{i}$ $=$ $\partial_{\mu}\hat{\lambda}^{i}+\hat{f}^{i}{}_{jk}\hat{\lambda}^{j}\hat{A}_{\mu}^{k}\,,$ (1.9)
$\delta_{\hat{\lambda}}\hat{B}_{\mu\nu}$ $=$ $-(\partial_{[\mu}\hat{\lambda}^{i})\hat{A}_{\nu]i}\,,$ (1.10)
$\delta_{\hat{\lambda}}\hat{g}_{\mu\nu}$ $=$ $0\,,$ (1.11)
where $\hat{\lambda}$ is a gauge parameter. Due to the previous restrictions, we will expand the gauge field following the proposal given in [22]
$\hat{A}_{\mu}^{i}$ $=c\ {\tau_{\mu}}^{-}\alpha_{-}^{i}+\frac{1}{c}a_{\mu}^{i}\,,$ (1.12)
where $v^{\pm}=\frac{1}{\sqrt{2}}(v^{0}\pm v^{1})$. The expansion (1.12) was constructed in such a way that the fundamental degrees of freedom of DFT remains convergent after taking the NR limit in the double geometry. In this way, we are sure that the NR heterotic supergravity can be rewritten in terms of $O(D,D+N)$ multiplets.
The main goal of this work is to explicitly prove that, when considering $c\rightarrow\infty$, the bosonic part of the heterotic Lagrangian,
$S_{\rm{het}}=\int d^{10}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}(\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}-\frac{1}{4}\hat{F}_{\mu\nu}{}^{i}\hat{F}^{\mu\nu}{}_{i})\,,$ (1.13)
with
$\bar{H}_{\mu\nu\rho}=3\left(\partial_{[\mu}\hat{B}_{\nu\rho]}-\hat{C}_{\mu\nu\rho}^{(g)}\right)\,,$ (1.14)
$\hat{C}_{\mu\nu\rho}^{(g)}=\hat{A}^{i}_{[\mu}\partial_{\nu}\hat{A}_{\rho]i}-\frac{1}{3}\hat{f}_{ijk}\hat{A}_{\mu}^{i}\hat{A}_{\nu}^{j}\hat{A}_{\rho}^{k}\,,$ (1.15)
$\hat{F}_{\mu\nu}{}^{i}=2\partial_{[\mu}\hat{A}_{\nu]}{}^{i}-\hat{f}^{i}{}_{jk}\hat{A}_{\mu}{}^{j}\hat{A}_{\nu}{}^{k}\,,$ (1.16)
remains finite. We will also present the explicit form of the gauge transformations and the heteorotic supergravity Lagrangian in its covariant form.
1.1 Organization of the paper
This paper is organized as follows: In the next section we review the cancellation of divergences in bosonic supergravity and we show the cancellation of divergencies in heterotic supergravity. In section 3 we study the gauge transformation rules of the fundamental fields of the theory upon considering the NR limit and we construct a suitable covariant derivative and curvatures for all the fields that transform in a non-covariant fashion. In section 4 we compute the finite Lagrangian of the theory. We will write not only the gauge part using new curvature for the gauge fields, but also the ungauged part of the Lagrangian [13] using new covariant structures which explicit show the invariance under infinitesimal diffeomorphisms. Finally, in section 5, we present a discussion and an outlook.
2 Cancellation of divergences
When one expands the action (1.7), both the $\hat{R}$ and $-\frac{1}{12}\hat{H}^{2}$ terms produce $c^{2}$ contributions. The same happens when we analyze (1.13), but this time the gauge compensation of $-\frac{1}{12}\bar{H}^{2}$ is given also by the $-\frac{1}{4}\hat{F}^{2}$ term. We will analyze both scenarios separately.
2.1 Ungauged case (review)
Let’s review the cancellation of divergences in the bosonic supergravity Lagrangian. To do so, we define
$\hat{R}^{\rho}{}_{\mu\nu\sigma}=-2\partial_{[\mu}\Gamma^{\rho}_{\nu]\sigma}-2\Gamma^{\rho}_{[\mu|\lambda}\Gamma^{\lambda}_{|\nu]\sigma}$ (2.1)
and $\hat{R}=\hat{R}^{\rho}{}_{\mu\nu\sigma}\delta_{\rho}^{\nu}\hat{g}^{\mu\sigma}$. The divergent part of the Ricci scalar, after imposing the Newton-Cartan relations (1.6), is given by
$\hat{R}^{(2)}=-\frac{c^{2}}{2}\eta_{ab}\partial_{\mu}{\tau_{\nu}{}^{a}}\partial_{\rho}{\tau_{\sigma}{}^{b}}(h^{\mu\sigma}h^{\nu\rho}-h^{\mu\rho}h^{\nu\sigma})\,.$ (2.2)
So far we have just considered the expansion of the metric tensor (1.1) and its inverse (1.2). Therefore, for backgrounds where $h^{\mu\nu}$ is constant, the divergence in the Ricci scalar becomes trivial and it is possible to define a non-relativistic supergravity without including a non-relativistic expansion in the $\hat{B}$-field. For a more general scenario we impose (1.3) and then the divergent part of $-\frac{1}{12}\hat{H}^{2}$ is given by
$-\frac{1}{12}(\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho})^{(2)}=-\frac{c^{2}}{2}\epsilon_{ab}\epsilon_{cd}\eta^{ac}\partial_{\mu}\tau_{\nu}{}^{b}\partial_{\rho}\tau_{\sigma}{}^{d}(h^{\mu\rho}h^{\nu\sigma}-h^{\mu\sigma}h^{\nu\rho})\,.$ (2.3)
Using the relation $\epsilon_{ab}\epsilon_{cd}\eta^{ac}=\eta_{bd}$ one can cancel the divergences coming from the Ricci scalar.
2.2 Gauged case
The incorporation of the gauge field $\hat{A}_{\mu}^{i}$ brings new divergences to the supergravity Lagrangian. On the one hand, the Yang-Mills Lagrangian,
$L_{\textrm{YM}}=-\frac{1}{4}\hat{F}_{\mu\nu}^{i}\hat{F}^{\mu\nu}_{i}$ (2.4)
upon considering the non-relativistic expansions (1.12) and (1.2) contains a divergence of the form (see appendix 5 for intermediate steps)
$L_{\textrm{YM}}^{(2)}=-\frac{c^{2}}{2}\alpha_{-}^{i}\alpha_{-i}\partial_{\mu}{\tau_{\nu}^{-}}\partial_{\rho}{\tau_{\sigma}^{-}}(h^{\mu\rho}h^{\nu\sigma}-h^{\mu\sigma}h^{\nu\rho})\,.$ (2.5)
These terms indicate that the $A$-expansion in (1.12) does not lead to a finite Yang-Mills theory. However, the inclusion of Chern-Simons terms in the heterotic Lagrangian (1.13) compensates the divergence (2.5). Indeed, one can compute (we use CADABRA [31] for performing this test. The details are given in the appendix),
$-\frac{1}{12}(\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho})^{(2)}=-R^{(2)}+\frac{1}{4}(\hat{F}_{\mu\nu}^{i}\hat{F}^{\mu\nu}_{i})^{(2)}\,.$ (2.6)
At this point, we have proved that the Heterotic supergravity Lagrangian is finite, but trying to write it in non-covariant form leads to huge number of terms. For this reason, our strategy will be to write it in covariant form, making use of the diffeomorphism and gauge invariance of the action.
In the next section, we will study the form of the gauge transformation rules and we will construct a suitable covariant derivative. After that we will construct new curvatures for all the fundamental fields transforming in a non-covariant way under gauge transformations.
3 Transformations and curvatures
In this section we will focus on two symmetries: infinitesimal diffeomorphism and the non-Abelian gauge transformations associated to the $A$-field.
3.1 Covariant derivative
The covariant derivative of $\tau_{\nu}{}^{a}$ can be easily computed,
$\nabla_{\mu}\tau_{\nu}{}^{a}$ $=$ $\partial_{\mu}\tau_{\nu}{}^{a}-\Gamma_{\mu\nu}^{\rho}\tau_{\rho}{}^{a}$
$=$ $\partial_{\mu}\tau_{\nu}{}^{a}+\partial_{(\mu}{\tau_{\nu)}{}^{a}}-\frac{1}{c^{2}}\partial_{(\mu}{\tau_{\nu)}{}^{a}}+T_{\mu\nu}^{a}\,,$
where $T_{\mu\nu}{}^{a}$ is given by the following covariant quantity,
$T_{\mu\nu}{}^{a}$ $=$ $\eta_{bc}\eta^{ad}\nabla_{(\mu|}{\tau_{\rho}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}-\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu|}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}$ (3.2)
$-\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{(\mu}{\tau_{\rho}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}+\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}\,.$
An interesting observation of $\nabla_{\mu}\tau_{\nu}{}^{a}$ is that this quantity does not diverge, but contains $\frac{1}{c^{2}}$ that we need to track in order to correctly write the finite Lagrangian. This means that the finite Lagrangian will evolve terms with powers of $c$, but this is just the price to pay for writing it in covariant form. The antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ satisfies
$\nabla_{[\mu}\tau_{\nu]}^{a}$ $=$ $\partial_{[\mu}\tau_{\nu]}{}^{a}$ (3.3)
while the symmetric part can be written as
$\partial_{(\mu}\tau_{\nu)}{}^{a}=\frac{c^{2}}{2c^{2}-1}\Big{(}\nabla_{(\mu}\tau_{\nu)}{}^{a}-T_{\mu\nu}{}^{a}\Big{)}=Q_{\mu\nu}{}^{a}\,.$ (3.4)
On the other hand the covariant derivative of $\tau^{\nu}{}_{a}$ is given by
$\nabla_{\mu}\tau^{\nu}{}_{a}=\partial_{\mu}\tau^{\nu}{}_{a}+\Gamma_{\mu\rho}^{\nu}\tau^{\rho}{}_{a}=\frac{1}{2}\partial_{\mu}\tau^{\nu}{}_{a}+S_{\mu}{}^{\nu}{}_{a}\,,$ (3.5)
where $S_{\mu}{}^{\nu}{}_{a}$ is a covariant quantity given by
$S_{\mu}{}^{\nu}{}_{a}$ $=$ $c^{2}\Big{(}\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\tau_{\mu}^{c}\tau^{\rho}_{a}h^{\nu\sigma}+\nabla_{[\mu}{\tau_{\rho]}^{b}}\eta_{ab}h^{\nu\rho}\Big{)}+\frac{1}{2}Q_{\mu\rho}{}^{b}(\tau^{\nu}_{b}\tau^{\rho}_{a}+\eta_{ab}h^{\nu\rho})$ (3.6)
$+\frac{1}{2}\nabla_{[\mu}{\tau_{\rho]}^{b}}(2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d}-\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}h^{\nu\rho})+\nabla_{[\rho}{\tau_{\sigma]}^{b}}(\eta_{bc}\tau_{\mu}^{c}\tau^{\sigma}_{a}h^{\nu\rho}+\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\rho}_{a}\tau^{\sigma}_{e})$
$+\frac{1}{4c^{2}}\Big{(}Q_{\mu\rho}{}^{b}(-2\tau^{\nu}_{b}\tau^{\rho}_{a}+2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+\nabla_{\mu}{\tau_{\rho}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})$
$-\nabla_{\rho}{\tau_{\mu}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+4\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\sigma}_{a}\tau^{\rho}_{e}\Big{)}\,.$
Considering the Newton-Cartan relation, it is easy to prove that the ungauged Lagrangian can be written in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$ and the tensors $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$, which also depends on the previous covariant derivatives. In other words, we do not need to compute the covariant derivative of $h_{\mu\nu}$ and $h^{\mu\nu}$ to write, for example, the Riemann tensor, the Ricci tensor or the Ricci scalar in covariant form.
3.2 Gauge transformations
Transformation rules under the non-Abelian gauge symmetry can be extracted from (1.9), (1.10) and (1.11). We will impose an extra expansion on the gauge parameter and the structure constants following [22],
$\hat{\lambda}^{i}$ $=$ $\frac{1}{c}\lambda^{i}$ (3.7)
$\hat{f}_{ijk}$ $=$ $cf_{ijk}\,.$ (3.8)
At this point, one can think that the rescaling of the structure constants might include divergences in the gauge Lagrangian. However, this reescaling is safe because higher order $\hat{f}_{ijk}$ contributions are $\propto\frac{1}{c}$. Then, the rescaling only introduces cubic contributions in gauge fields in the finite Lagrangian, which we will present in the next section.
After exploring the gauge transformation (1.9), (1.10) and (1.11) we find the following rules upon taking the limit $c\rightarrow\infty$,
$\delta_{\lambda}a_{\mu}^{i}$ $=$ $\partial_{\mu}\lambda^{i}+f^{i}{}_{jk}\lambda^{j}a_{\mu}^{k}\,,$ (3.9)
$\delta_{\lambda}\alpha_{-}^{i}$ $=$ $f^{i}{}_{jk}\lambda^{j}\alpha_{-}^{k}\,,$ (3.10)
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-(\partial_{[\mu}\lambda^{i})\alpha_{-i}\tau_{\nu]}^{-}\,,$ (3.11)
$\delta_{\lambda}h_{\mu\nu}$ $=$ $0\,,$ (3.12)
$\delta_{\lambda}h^{\mu\nu}$ $=$ $0\,,$ (3.13)
$\delta_{\lambda}\tau_{\mu}{}^{a}$ $=$ $0\,,$ (3.14)
$\delta_{\lambda}\tau^{\mu}{}_{a}$ $=$ $0\,.$ (3.15)
From the previous transformations, we learn that $a_{\mu}^{i}$ transforms as a gauge connection, $\alpha_{-}^{i}$ transforms as a vector and $b_{\mu\nu}$ transforms in a non-covariant way, emulating a Green-Schwarz mechanism. Interestingly, this mechanism cannot be eliminated through field redefinitions in the relativistic scenario. However, when one considers the NR limit, it is possible to define the field
$\bar{b}_{\mu\nu}=b_{\mu\nu}+a_{[\mu|i}\alpha_{-}^{i}\tau_{|\nu]-}$ (3.16)
which is gauge invariant, i.e. $\delta_{\lambda}\bar{b}_{\mu\nu}=0$. This is an interesting property of the heterotic supergravity under the NR limit which will help us to construct a suitable curvature for the $\bar{b}$-field.
3.3 Gauging the covariant derivative
Considering an arbitrary gauge vector $v^{i}$ upon taking the NR limit, the partial derivative acting on this vector is not covariant. Therefore we need to define a covariant derivative for the gauge symmetry in the following way,
$\nabla_{\mu}v^{i}=\partial_{\mu}v^{i}-f^{i}{}_{jk}a_{\mu}{}^{j}v^{k}\,.$ (3.17)
We choose $a_{\mu}{}^{j}$ over $\alpha_{-}^{i}$ to construct the covariant derivative due to the transformation property of $a_{\mu}{}^{j}$. In this case, we need a proper connection and $\alpha_{-}^{i}$ transforms as a gauge vector.
Let’s start constructing the curvature for the gauge connection $a_{\mu}^{i}$, which is given by
$f_{\mu\nu}{}^{i}=2\partial_{[\mu}a_{\nu]}{}^{i}-f^{i}{}_{jk}a_{\mu}{}^{j}a_{\nu}{}^{k}\,,$ (3.18)
The field $\alpha_{-}^{i}$ is already covariant, so we do not need to construct its curvature. If we consider $\bar{b}$ as the fundamental field of the theory, then we define its curvature as
$\bar{h}_{\mu\nu\rho}=3\partial_{[\mu}\bar{b}_{\nu\rho]}\,=h_{\mu\nu\rho}-3c^{(g)}_{\mu\nu\rho}.$ (3.19)
From the previous curvature, we can identify the equivalent quantity to the Chern-Simons terms in this limit,
$c^{(g)}_{\mu\nu\rho}=-\partial_{[\mu}(a_{\nu}^{i}\tau_{\rho]}^{-}\alpha_{-i}).$ (3.20)
In the next section, we will use these curvatures to present the finite NR gauge Lagrangian.
4 Finite Lagrangian
In this section we construct all the finite contributions to the heterotic supergravity Lagrangian in covariant form upon taking the NR limit. For the readers interested only in the final result, we report the final form of the gauge Lagrangian in the conclusions.
4.1 Ungauged Lagrangian
We would like to present the finite heterotic Lagrangian in terms of covariant quantities. We will choose objects which are covariant under infinitesimal diffeomorphisms. Due to the previous section, we have all the pieces to write the ungauged part of the finite Lagrangian in its covariant form. This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term and the H-term given by
$L$ $=$ $R^{(0)}+4\partial_{\mu}\varphi\partial^{\mu}\varphi-\frac{1}{12}H^{(0)}_{\mu\nu\rho}H^{(0)}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$ (4.1)
$-\frac{1}{2}H^{(0)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}H^{(2)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\eta^{cd}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c}\tau^{\sigma}{}_{b}\tau^{\gamma}{}_{d}h^{\rho\epsilon}\,,$
where
$H^{(0)}_{\mu\nu\rho}$ $=$ $h_{\mu\nu\rho}=3\partial_{[\mu}b_{\nu\rho]}\,,$ (4.2)
$H^{(2)}_{\mu\nu\rho}$ $=$ $-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,,$ (4.3)
and the Ricci scalar is given by
$R^{(0)}$ $=$ $8\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau^{\mu}{}_{a}}h^{\nu\rho}+5\nabla_{[\mu}{\tau_{\nu]}{}^{c}}\eta^{ab}(\nabla_{\rho}{\tau^{\mu}{}_{a}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{b}+\nabla_{\rho}{\tau^{\rho}{}_{a}}\tau^{\mu}{}_{b}\tau^{\nu}{}_{c})$ (4.4)
$-\frac{3}{8}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}(\tau^{\nu}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\rho}-2\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\rho}+\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma})$
$+\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\rho}h^{\nu\sigma}+4c^{2}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\sigma}h^{\nu\rho}$
$+\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}-\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma}-\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho})$
$+\frac{1}{c^{2}}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho}-2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma})$
$-2\nabla_{\mu}{\nabla_{\nu}{\tau_{\rho}{}^{a}}}\tau^{[\rho}{}_{a}h^{\nu]\mu}+f(Q,S)\,,$
where $f(Q,S)$ is a covariant contribution since both $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$ are covariant,
$f(Q,S)$ $=$ $8Q_{\mu\nu}{}^{a}(S_{\rho}{}^{[\mu}{}_{a}h^{\rho]\nu}+\nabla_{\rho}{\tau^{[\rho}{}_{a}}h^{\mu]\nu})-8\nabla_{[\nu}{\tau_{\rho]}{}^{a}}S_{\mu}{}^{\nu}{}_{a}h^{\mu\rho}-\frac{1}{2}Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma}$ (4.5)
$+Q_{\mu\nu}{}^{a}\eta^{bc}(-2\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}+\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}+\nabla_{\rho}{\tau^{\rho}{}_{b}}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
$+Q_{\mu\nu}{}^{a}\eta^{bc}(2S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}-S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}-S_{\rho}{}^{\rho}{}_{b}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
$+\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}(-\frac{3}{2}\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+2\tau^{\mu}{}_{a}\tau^{[\rho}{}_{b}h^{\sigma]\nu}+2\tau^{\rho}{}_{b}\tau^{\sigma}{}_{a}h^{\mu\nu}+4\tau^{\mu}{}_{b}\tau^{[\rho}{}_{a}h^{\sigma]\nu})$
$+\nabla_{\nu}{\tau_{\rho}{}^{c}}\eta^{ab}(-\frac{5}{2}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+5S_{\mu}{}^{\mu}{}_{a}\tau^{[\nu}{}_{c}\tau^{\rho]}{}_{b})$
$+\frac{5}{2}\nabla_{\rho}{\tau_{\nu}{}^{c}}\eta^{ab}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}h^{\mu[\nu}h^{\rho]\sigma}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\rho}{}_{d}h^{\nu]\sigma}$
$+3\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\sigma}{}_{d}h^{\rho]\nu})+\frac{4}{c^{2}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}Q_{\mu\nu}{}^{a}(\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+\tau^{\mu}{}_{b}\tau^{\sigma}{}_{a}h^{\nu\rho})$
$+4\nabla_{\mu}{Q_{\nu\rho}{}^{a}}\tau^{[\mu}{}_{a}h^{\nu]\rho}\,.$
In $R^{(0)}$ we can explicitly see that terms proportional to $c^{2}$ and $\frac{1}{c^{2}}$ naturally arises when we write this object using covariant derivatives. When we replace the covariant derivatives by their definitions, all these divergent and convergent quantities cancel, giving a finite contribution in terms of partial derivatives.
4.2 Gauged terms
Before analyzing the gauge contributions of the heteorotic Lagrangian, let’s focus on the ungauged terms depending on the b-field,
$L_{b}$ $=$ $-\frac{1}{4}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\rho}h^{\nu\gamma}+\frac{1}{2}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}$ (4.6)
$-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\rho}{}_{c}h^{\mu\sigma}h^{\nu\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$
These terms are not gauge invariant, due to the transformation (3.11). Therefore, we rewrite these contributions in terms of the new curvature for the b-field,
$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}+\nabla_{\mu}\tau_{\nu}^{-}f_{\gamma\rho}{}^{i}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+L_{\textrm{comp}}\,,$ (4.7)
where $L_{\textrm{comp}}=L^{(2)}_{\textrm{non-cov}}+L^{(3)}_{\textrm{non-cov}}+L^{(4)}_{\textrm{non-cov}}$ with
$L^{(2)}_{\textrm{non-cov}}$ $=$ $\nabla_{\mu}\tau_{\nu}^{-}a_{[\rho}^{i}\nabla_{\gamma]}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+3\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}a_{[\sigma}^{i}\alpha_{-i}\nabla_{\rho}\tau_{\gamma]}^{-}\tau^{\sigma}_{c}h^{\mu\gamma}h^{\nu\rho}$ (4.8)
$+\frac{1}{2}h_{\mu\nu\rho}a_{\gamma}^{i}\alpha_{-i}\nabla_{\sigma}(\tau_{\epsilon}^{-})h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,,$
$L^{(3)}_{\textrm{non-cov}}=-\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a_{\gamma}^{j}a_{\rho}{}^{k}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}\,,$ (4.9)
and
$L^{(4)}_{\textrm{non-cov}}$ $=$ $\frac{3}{4}a_{[\nu}^{i}\alpha_{-i}\nabla_{\mu}(\tau_{\rho]^{-}})a_{\gamma}^{j}\alpha_{-j}\nabla_{\sigma}\tau_{\epsilon}^{-}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,.$ (4.10)
In the next subsection we will study the contributions to the gauge Lagrangian coming from the Yang-Mills terms and the Chern-Simons contributions. We split the presentation according to an expansion in gauge fields.
4.3 Quadratic gauged contributions
4.3.1 Quadratic Yang-Mills contributions
The quadratic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
$+2\alpha_{-i}\eta^{ab}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau_{\sigma}^{-}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-2\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}\nabla_{[\rho}a_{\sigma]}^{i}h^{\mu\rho}h^{\nu\sigma}\,.$ (4.11)
The previous expression can be written in a covariant form as
$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}f_{\rho\sigma}^{i}h^{\mu\rho}h^{\nu\sigma}+L^{(3)}(a,\alpha)$ (4.12)
where the inclusion of the (gauge) covariant derivative and the curvature $f_{\mu\nu i}$ introduces cubic contributions
$L^{(3)}(a,\alpha)=-\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a^{j}_{\rho}a^{k}_{\sigma}h^{\mu\rho}h^{\nu\sigma}+2\alpha_{-i}\eta^{ab}f^{i}{}_{jk}a_{\mu}^{j}\alpha_{-}^{k}\nabla_{[\nu}\tau_{\rho]}^{-}\tau_{\sigma}^{-}\tau_{a}^{\nu}\tau_{b}^{\sigma}h^{\mu\rho}\,.$ (4.13)
4.3.2 Quadratic Chern-Simons contributions
The quadratic contributions coming from $-\frac{1}{12}\bar{H}^{2}$ are given by
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}a_{\sigma]i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}h_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}\,.$ (4.14)
Among these terms we have identified all the quadratic non-covariant contributions needed in (4.7). We can write the previous expression in covariant form as
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}+\tilde{L}^{(3,4)}(a,\alpha)\,,$ (4.15)
where once again the inclusion of the curvatures $f_{\mu\nu i}$ and $\bar{h}_{\rho\sigma\gamma}$ in the last line of the previous expression introduces cubic and quartic contributions, respectively, that we will take into account when we analyze the next orders. We capture all these new contributions under the notation $\tilde{L}^{(3,4)}(a,\alpha)$, which splits in the following form
$\tilde{L}^{(3)}(a,\alpha)=-\alpha_{-}^{i}\nabla_{\mu}\tau_{\nu}^{-}f_{i}{}^{jk}a_{\rho j}a_{\sigma k}h^{\mu\sigma}h^{\nu\rho}\,,$ (4.16)
and
$\tilde{L}^{(4)}(a,\alpha)=\frac{3}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}a_{[\sigma}^{j}\alpha_{-j}\nabla_{\rho}(\tau_{\gamma]}^{-})\tau_{+}^{\rho}h^{\mu\sigma}h^{\nu\gamma}\,.$ (4.17)
4.4 Cubic gauge contributions
4.4.1 Cubic Yang-Mills contributions
The cubic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
$-L^{(3)}_{\textrm{non-cov}}-L^{(3)}(a,\alpha)-\tilde{L}^{(3)}(a,\alpha)\,.$ (4.18)
The previous expression exactly cancels the contributions obtained in (4.13) and (4.16) and there are no additional contributions at this order in fields.
4.5 Quartic gauge contributions
4.5.1 Quartic Chern-Simons contributions
The quartic terms in $-\frac{1}{12}\bar{H}^{2}$ are given by
$-\tilde{L}^{(4)}(a,\alpha)\,,$ (4.19)
and there are no other additional contributions at this order, or higher, in fields.
5 Discussion
In this work we have studied the finite form of the heterotic supergravity Lagrangian, extending the formulation of [13]. The Ricci scalar part of the Lagrangian has been rewritten in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$, $Q_{\mu\nu}{}^{a}$ (eq. (3.4)) and $S_{\mu}{}^{\nu}{}_{a}$ (eq. (3.6)), showing the explicit diffeomorphism invariance of the theory.
When we include the heteoric gauge field, the bosonic part of the Lagrangian of this theory contains two divergent parts given by the Chern-Simons terms and the Yang-Mills Lagrangian which exactly compensate each other, as it happens in bosonic supergravity, giving rise to a finite Lagrangian. The gauge Lagrangian depends on a pair of gauge fields, $a_{\mu}^{i}$ and $\alpha_{-}^{i}$, the former transforming as a gauge connection, so we constructed its curvature, and the latter transforming as a gauge vector, so we do not need to construct a curvature. Interestingly enough, the Green-Schwarz mechanism for the b-field can be trivialized in this limit and therefore one can define $\tilde{b}_{\mu\nu}$ as in (3.16), and its curvature $\tilde{h}_{\mu\nu\rho}$, both of them gauge invariant. The gauge contributions in $\tilde{h}_{\mu\nu\rho}$ can be though of as Chern-Simons contributions.
In this work we have presented the finite Lagrangian of the theory which contains quadratic, cubic and quartic contributions. The full gauge Lagrangian is given by
$L^{(g)}_{\textrm{het}}$ $=$ $-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$ (5.1)
$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}$
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}$
$-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}\,,$
while the ungauged action must include the Chern-Simons terms in the curvature of the b-field as we show in (4.7), i.e.,
$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$ (5.2)
5.1 Outlook
As promising avenue for future work, it would be very interesting to explore the incorporation of $\alpha´$-corrections to both the bosonic and the heterotic supergravity  [^2] See [32] for a pedagogical review on this topic and references therein. in the NR limit. Since the Riemann tensor contains divergences, one should study whether these are controlled by the $\hat{H}_{\mu\nu\rho}$ (or $\bar{H}_{\mu\nu\rho}$ in the heterotic case) contributions, as happens at the two derivative level. The covariant way of writing the Ricci scalar using the covariant derivative of $\tau_{\mu}{}^{a}$ and $\tau^{\mu}{}_{a}$ could be easily extended in order to incorporate the spin connection $\omega_{\mu ab}$, so that the action is also invariant under the $SO(1,1)$ transformations. In this way, the higher-derivative contributions could be written in a more compact form, since we expect large contributions if the curvatures are not chosen in an optimal way.
While DFT has been proved a powerful tool to read the expansion of the A-field [22], taking the NR limit at the DFT level including higher-derivative terms is probably the best strategy to read these corrections and then parametrize them to the supergravity level.
Acknowledgements
E.L. is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.
Appendix A Details on the cancellation of divergences: gauge case
The field strength $F_{\mu\nu}{}^{i}$, after expanding $A_{\mu}{}^{i}$ considering (1.12) is given by
$F^{(1)}_{\mu\nu}{}^{i}=\alpha_{-}^{i}\partial_{\mu}{\tau_{\nu}^{-}}+\partial_{\mu}{\alpha_{-}^{i}}\tau_{\nu}^{-}-\alpha_{-}^{i}\partial_{\nu}{\tau_{\mu}^{-}}-\partial_{\nu}{\alpha_{-}^{i}}\tau_{\mu}^{-}\,,$ (A.1)
$F^{(0)}_{\mu\nu}{}^{i}=+\alpha_{-}^{j}\eta^{ik}\tau_{\mu}^{-}a_{\nu}^{i1}f_{jki1}-\alpha_{-}^{j}\eta^{ik}\tau_{\nu}^{-}a_{\mu}^{i1}f_{jki1}\,,$ (A.2)
$F^{(-1)}_{\mu\nu}{}^{i}=\partial_{\mu}{a_{\nu}^{i}}-\partial_{\nu}{a_{\mu}^{i}}\,,$ (A.3)
$F^{(-2)}_{\mu\nu}{}^{i}=-\eta^{ij}a_{\mu}^{k}a_{\nu}^{i1}f_{jki1}\,,$ (A.4)
where the notation $F^{(i)}_{\mu\nu}{}^{i}$ denotes the terms of order $c^{i}$ in the expansion. The contraction of this field strength with two inverse metrics, $F^{\mu\nu i}=F_{\rho\sigma}g^{\mu\rho}g^{\nu\sigma}$ yields the expression,
$F^{\mu\nu i}=F^{(1)\mu\nu i}+\mathcal{O}\big{(}\frac{1}{c}\big{)}\,,$ (A.5)
with
$F^{(1)\mu\nu i}=\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\sigma}h^{\nu\rho}\,.$ (A.6)
As the reader can easily check, the only contraction which produces non-vanishing divergences is given by $F_{(1)\mu\nu}{}^{i}F^{(1)\mu\nu j}\eta_{ij}$, which leads to the result (2.5).
For the $-\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}$ contributions, an important observation is that
$\bar{H}_{\mu\nu\rho}=\bar{H}^{(2)}_{\mu\nu\rho}+\bar{H}_{(0)\mu\nu\rho}+\mathcal{O}\big{(}\frac{1}{c}\big{)}\,,$ (A.7)
while the same object when all the indices raised by the inverse metric has the following form,
$\bar{H}^{\mu\nu\rho}=\bar{H}^{(0)\mu\nu\rho}+\mathcal{O}\big{(}\frac{1}{c^{2}}\big{)}\,.$ (A.8)
Thus, we focus on the objects that contribute to divergences,
$\bar{H}^{(2)}_{\mu\nu\rho}$ $=$ $-\epsilon_{ab}\partial_{\mu}{\tau_{\nu}^{a}}\tau_{\rho}^{b}+\epsilon_{ab}\partial_{\mu}{\tau_{\rho}^{a}}\tau_{\nu}^{b}-\epsilon_{ab}\partial_{\nu}{\tau_{\rho}^{a}}\tau_{\mu}^{b}+\epsilon_{ab}\partial_{\nu}{\tau_{\mu}^{a}}\tau_{\rho}^{b}-\epsilon_{ab}\partial_{\rho}{\tau_{\mu}^{a}}\tau_{\nu}^{b}$ (A.9)
$+\epsilon_{ab}\partial_{\rho}{\tau_{\nu}^{a}}\tau_{\mu}^{b}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\nu}{\tau_{\rho}^{-}}\tau_{\mu}^{-}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\rho}{\tau_{\mu}^{-}}\tau_{\nu}^{-}$
$-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{-}}\tau_{\rho}^{-}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\rho}{\tau_{\nu}^{-}}\tau_{\mu}^{-}$
$+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\mu}{\tau_{\rho}^{-}}\tau_{\nu}^{-}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\nu}{\tau_{\mu}^{-}}\tau_{\rho}^{-}\,,$
$\bar{H}^{(0)\mu\nu\rho}$ $=$ $h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\rho}_{c}h^{\mu\sigma}h^{\nu\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\nu}_{c}h^{\mu\sigma}h^{\rho\gamma}$ (A.10)
$+\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\mu}_{c}h^{\nu\sigma}h^{\rho\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\rho}_{c}h^{\mu\gamma}h^{\nu\sigma}+\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\nu}_{c}h^{\mu\gamma}h^{\rho\sigma}$
$-\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\mu}_{c}h^{\nu\gamma}h^{\rho\sigma}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\mu}_{a}\tau^{\epsilon}_{b}h^{\nu\sigma}h^{\rho\gamma}$
$-\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\nu}_{a}\tau^{\epsilon}_{b}h^{\mu\gamma}h^{\rho\sigma}$
$-\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\gamma}h^{\nu\epsilon}h^{\rho\sigma}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\rho}_{a}\tau^{\epsilon}_{b}h^{\mu\sigma}h^{\nu\gamma}$
$-\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\mu}_{a}\tau^{\epsilon}_{b}h^{\nu\gamma}h^{\rho\sigma}$
$+\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\epsilon}h^{\nu\gamma}h^{\rho\sigma}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\nu}_{a}\tau^{\epsilon}_{b}h^{\mu\sigma}h^{\rho\gamma}$
$+\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\sigma}h^{\nu\epsilon}h^{\rho\gamma}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\rho}_{a}\tau^{\epsilon}_{b}h^{\mu\gamma}h^{\nu\sigma}$
$+\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}\,.$
The multiplication of the previous two quantities leads directly to the result (2.6).